find the equation of a parabola with a focus at (0,-1) and a directrix at y=4
1
Expert's answer
2020-12-15T01:24:25-0500

"\sqrt{(x-0)^2+(y+1)^2}=y-4\\
\text{square both sides}\\
(x-0)^2+(y+1)^2=(y-4)^2\\
x^2+y^2+2y+1=y^2-8y+16\\
x^2+y^2-y^2+2y+8y+1-16=0\\
x^2+10y-15=0"